OpenStudy (anonymous):

how do you solve 2(x-3)^2=16

OpenStudy (anonymous):

PEMDAS

OpenStudy (anonymous):

2(x-3)(x-3)=16 divide by two,both sides, FOIL, then continue

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

you could also use a calculator graph y=16 then y=2(x-3)^2 and find the intersection

OpenStudy (anonymous):

\[2(x-3)^2= 16 \Rightarrow (x-3)^2 = 8 \Rightarrow x-3 = \pm \sqrt{8} \Rightarrow x = 3\pm \sqrt{8}\]